OpenStudy (anonymous):

in an examination if a student looses mark for the wrong answer and get 1 mark for correct answer if he got 30 mark for solving 54 que. then how many que. were correct?

OpenStudy (ghazi):

there could be many possibilities of getting 30 outta 54

OpenStudy (unklerhaukus):

R-W=30 R+W=54 add these equations together 2R=... R=... R+W=54 W=...

OpenStudy (unklerhaukus):

ghazi i think 54 was the number of questions attempted , the examination probably had more questions than this

OpenStudy (ghazi):

you sure... :) ? that's why i wrote many possibilities also what if marks deducted for one wrong answer is not exactly 1 , what is for one wrong question it is 0.25 or 0.50

OpenStudy (ghazi):

*if

OpenStudy (anonymous):

The last thing that @rajmscrewa wrote on the previous post was that each wrong answer was one point.

OpenStudy (anonymous):

ok solve all possibilities

OpenStudy (anonymous):

@UnkleRhaukus gave a good method for solving above. It leads to one possibility.

OpenStudy (ghazi):

yea agreed...for one negative..it's just one answer

OpenStudy (anonymous):

ok if he looses 0.50 marks for a wrong answer than

OpenStudy (anonymous):

I thought about that a little already... I think that it would start: R-0.5W=30 R+W=54 Where R is number correct and W is number incorrect.

OpenStudy (anonymous):

than

OpenStudy (anonymous):

if he looses 0.25 marks

OpenStudy (anonymous):

I'm not sure I'm right... it doesn't seem to work out.

OpenStudy (unklerhaukus):

' student looses mark ' this is singular tense even if there is a word missing

OpenStudy (anonymous):

Wait, yeah it worked. It becomes formulaic. (weight for correct)R-(weight for incorrect)W=score R+W=(number of questions)